Wang Supply prepares monthly financial statements. Listed below are some selected accounts and their balances at June 30. Adjustments generally are done monthly, but no adjustments have been made for the month of June, 2017. Wang has a May 31 year end.

Wang Supply
Unadjusted Trial Balance (partial)
June 30, 2017

Debit Credit
Prepaid Insurance $ 2,940
Supplies 520
Vehicles 32,000
Accumulated Depreciation - Vehicles $ 9,360
Notes Payable 22,400
Unearned Revenue 1,500
Rent Revenue 142,000

An analysis of the account balances by the company's bookkeeper provided the following additional information:

1. A two-year insurance policy was purchased on March 1, 2017 for $3,360.
2. A physical count of supplies revealed $250 on hand on June 30, 2017.
3. The note payable is for one year at an annual interest rate of 5%, dated June 1, 2017.
4. The vehicle is being depreciated at the rate of $2,520 per year.
5. The amount of rent received in advance that remains unearned at June 30, 2017 is $1,370.

Using the additional information above, prepare the adjusting entries that should be made by the Wang Supply on June 30. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Round answers to 0 decimal places, e.g. 125.)
Answer #1
DATE ACCOUNT TITLE DEBIT CREDIT
June 30,2017
1 Insurance expense 140
Prepaid Insurance 140

[Being insurance expired for 4 months out of 2 years or 24 months (2years*12 months per year).

Insurance expense per month :3360/24 = $ 140 per month

2 Supplies expense 270
Supplies 270
[Being supplies used :520 Unadjusted balance -250 at end]
3 Interest expense 93
Interest payable 93
[Being interest accrued for month (1June -30June) :22400*.05*1/12 = 93.33 ]
4 Depreciation expense 210
Accumulated Depreciation - Vehicles 210
[Being depreciation recorded for one month :2520/12= 210]
5 Unearned revenue 130
Rent revenue 130
[Being rent revenue earned :1500 Unadjusted balance -1370 Balance at end ]